PLEASE TYPE:) thank you (show the work for math)

You are a pediatric nurse taking care of Takesha, a 3 year-old, 14.2-kg, patient on the medical floor. Takesha was admitted to the hospital last evening with and an acute asthma exacerbation and hypoxia. She is on oxygen at 2 liters/min and is maintaining oxygen sats around 95%. She is receiving albuterol treatments every 4 hours and will start prednisolone this morning. She has a peripheral IV of D5W .25NS + 20 mEq KCl/liter at standard rate. This morning you received the following orders;

  • Prednisilone 15mg/5ml; 4 ml q12h
  • Convert IV to saline trap if adequate po intake
  • Wean oxygen to maintain sats > 93%

            You walk into Takesha’s room to complete her morning assessment. The mother is sitting in the chair beside the bed.

Questions:

1. Is Takesha’s prednisolone dose (above) appropriate based on the recommended safe dose of 1-2mg/kg/day?

2. Calculate Takesha’s recommended IV rate/daily fluid requirement based on the standard formula below.

100ml/kg for 1st 10 kg of body weight

50ml/kg for next 10kg (between 10-20 kg) of body weight

20ml/kg for any additional body weight in kg.

Total amount/divide by 24 for hourly rate.

Total fluid requirement:                        Hourly rate:

3a. What is Takesha’s developmental stage?

3b. What are some nursing considerations needed when caring for her?

4. Takesha will be going home in a few days. What are some anticipatory guidance/safety concerns r/t her age/developmental stage you need to address before discharge?
Answer #1

1. Prednisolone:

Safe dose= 1-2mg/kg/day

= 2×14.2/day

=28.4mg/day

Available: 15mg/5ml

3mg/1ml

Child dose: 4ml

That is

12mg, q12hour

So, 24mg/day

Upto 38mg/day safe dose,

So this 24mg/day is safe dose.

2. IV fluid,

100ml/kg First 10kg

So, 1000ml

50ml/kg 10-20kg

50×4.2= 210ml

Here, child weight=10+4.2kg=14.2kg

Fluids=1000ml+210ml=1210ml.

Total fluid required = 1210ml

For hourly= 1210/24=50ml

3a, Development stage: Autonomy versus Shame.

3b,

Strict intake output chart

Monitor blood sugar closely.

Infection prevention.

Adequate hydration.

Aseptic precautions.

Monitor electrolytes level carefully.

4.

Give safe environment.

Electrical safety

Safety gates for stairs.

Avoid over polishing floors.